Tags

Fast Luminosity Monitor Electronic Logbook 2005

23 Jul 2024 - 17:38

Hardware Layout

June 2005

June 07, 2005 (E. Tanke): FLM Channel Map

V1 V2 V3 V4 V5 V6 V7 V8
FLM INT CUR: 1 2 3 4 5 6 7 8 (Cabling Order)
DIGI_CHAN_MAP: 7 6 5 4 3 2 1 0 (TSharc Module Input Board)
VERT_CHAN_MAP: 0 2 4 6 1 3 5 0 (Geometric Order)

VERT_CHAN_MAP runs from bottom to top of detector where 0 means not connected.

H1 H2 H3 H4 H5 H6 H7 H8 H9 H10 H11 H12 H13 H14 H15 H16
FLM INT CUR: 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 (Cables)
DIGI_CHAN_MAP: 23 22 21 20 19 18 17 16 15 14 13 12 11 10 9 8 (TSharc)
HORZ_CHAN_MAP: 13 9 5 1 0 11 7 3 2 6 10 14 4 8 12 0 (Geom)

Note: 14-Sep-2005 --> Corrected the H DIGI_CHAN_MAP; the higher and lower 8 H channels were inverted.

-- EugeneTanke - 07 Jun 2005

June 10, 2005 (E. Tanke): FLM Bunch Map

Map of DSP bunches [0..182] to CESR Train and Bunch
Bunch 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21
T9 130 131 132 133 134 135 136 137 138 139 140 141 142 143 144 145 146 147 148 149 150
T1 151 152 153 154 155 156 157 158 159 160 161 162 163 164 165 166 167 168 169 170  
T2 171 172 173 174 175 176 177 178 179 180 181 182 0 1 2 3 4 5 6 7  
T3 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25 26 27 28
T4 29 30 31 32 33 34 35 36 37 38 39 40 41 42 43 44 45 46 47 48  
T5 49 50 51 52 53 54 55 56 57 58 59 60 61 62 63 64 65 66 67 68  
T6 69 70 71 72 73 74 75 76 77 78 79 80 81 82 83 84 85 86 87 88 89
T7 90 91 92 93 94 95 96 97 98 99 100 101 102 103 104 105 106 107 108 109  
T8 110 111 112 113 114 115 116 117 118 119 120 121 122 123 124 125 126 127 128 129  
-- EugeneTanke - 10 Jun 2005

June 26, 2005 (M.Palmer, E. Tanke): Re-installed the FLM hardware in the tunnel at 3W

  • Re-installed the FLM detector hardware in the tunnel at 3W
  • The following notes are appended from prep work last week
    • On June 20, RJY attempted to install the FLM HV ready-chain and power supplies but immediately encountered ready-chain problems
      • Subsequent tests with GWC/WGD/JJB indicated a problem in the HV comparitor chassis
      • System was removed from tunnel
    • On June 21, further checkout was carried out by RJY/JJB
      • Ultimately found that an improperly grounded low voltage power supply in the system
      • Bench tests worked after grounding problem fixed
    • On June 22, RJY successfully re-installed the HV system in the tunnel.
-- EugeneTanke - 26 Jun 2005

June 27, 2005 (M. Palmer, E. Tanke): Checkout of data acquired on June 26

  • Have discovered problem with "analog readout" of FLM which was causing geometric distortions during tests yesterday evening
    • Prior to introduction of the FLM Digital Electronics, pedestals were obtained by setting pedestal values in the DB to 0 and gains to 1, and then monitoring channel by channel signal values in the FLM AVG CUR node which then contained simple averages of the values from the FLM INT CUR node for which the pedestals are needed.
    • With the introduction of the digital readout and the noisier environment seen by the analog readout, the FLM AVG CUR now contains "corrected" signals where the readout from the unused channels on each card is used to removed the common mode noise introduced in the digital box. This means that the above pedestal measurement scheme is no longer valid!!! It introduces a channel-to-channel offset which is fairly negligible at high photon rates but not at the rates we are presently at.
    • Have replaced the former routine that handled pedestal measurements [cesr.flm.do_gains]load_gains.f77 with a new one that does only pedestals based on direct averaging of the FLM INT CUR node. The new program is in [cesr.flm.do_gains]load_peds.f77 . It no longer modifies the pedestal and gain values in the database, but simply collects FLM INT CUR pedestal values and makes them available for loading into the FLM PED CUR node.
  • Have also reviewed the timing scan and multi-channel analyzer output obtained yesterday evening. These appear to be consistent with proper FLM digital operation. We will need to take some better statistics data tonight to precisely quantify performance.
-- MarkPalmer - 27 Jun 2005

June 28, 2005 (M. Palmer, E. Tanke): FLM measurements

  • FLM measurements were made under low luminosity conditions; the measurement files are in the subdirectories of tanke/cesr/ts101/flm03w/ :
    • timing scan with test_daqpavg:
      • test_daqpavg/2005_06_28/FLM_PAW_2005_06_28_0200.DAT
    • MCA data (NBIN=250, MCAMIN= -7500, MCAMAX=500, NLOOPS=10, timing=1380) :
      • test_24c_mca/2005_06_28/FLM_PAW_MCA_2005_06_28_0220.DAT
      • test_24c_mca/2005_06_28/FLM_PAW_NBIN_2005_06_28_0220.DAT
-- EugeneTanke - 28 Jun 2005

June 29, 2005 (M. Palmer, E. Tanke, J. Hylas, T. Banta) FLM Digital Checkout and Emittance Measurements

  • Lost nearly half of our time to power glitch (lightening) and recovery
  • 0430-0530 JTH tuned up some acceptable colliding beam conditions for testing
  • Standard FLM readout nodes being logged with FLM_DLOG into data set 31
  • FLM Digital Readout Tests:
    • Timing scan:
      • test_daqpavg/2005_06_29/FLM_PAW_2005_06_29_0515.dat
      • Peak of timing appears to be around 1440 (old timing used to be about 1340)
    • MCA data (NBIN=250, MCAMIN=-7500, MCAMAX=500, NLOOPS=100):
      • Initially use channels 1-6
      • test_24c_mca/2005_06_29/0535/FLM_PAW_MCA_2005_06_29_0535.DAT
      • test_24c_mca/2005_06_29/0535/FLM_PAW_NBIN_2005_06_29_0535.DAT
      • See some signal
    • MCA data (NBIN=250, MCAMIN=-7500, MCAMAX=500, NLOOPS=250):
      • Uses ALL channels
      • Approximately 15 minute data acquisition
      • test_24c_mca/2005_06_29/0550/FLM_PAW_MCA_2005_06_29_0550.DAT
      • test_24c_mca/2005_06_29/0550/FLM_PAW_NBIN_2005_06_29_0550.DAT
    • 0600 - Check digitization impact on analog signal by cycling digital data acquisition using the run_adc com file
      • Toggles acquisition approx. every 5 seconds
      • See approx. 800 units modulation in the FLM VRAW signal
    • 0625 - (06:24:54 to be precise) Get continuous bunch acquisition running
      • With limited duty cycle don't observe a significant shift in analog readout
      • test_6c_cont program settings: Luminosity turns=10000, individual channel turns=250, threshold=-1200, timing=1440
      • Stop at 0635; files:
        • test_6c_cont/2005_06_29/0625/flm_paw_ind_2005_06_29_0625.dat
        • test_6c_cont/2005_06_29/0625/flm_paw_lum_2005_06_29_0625.dat
      • The lum file appears reasonable, however, there is an apparent problem with the individual channel geometric file (see note at 0725)
    • 0643 - 0645 Check response versus VBLIMING 2
    • Current Dependent Emittance Checks
      • 0658 1.5 mA e+ against 0.25 mA e-
        • Have B1 lifetime problems
        • Save with QTUNEING 6 1777 --> 1769 and XQUNEING 4 -58 --> -61
      • 0703-0710 Refill and take data 55/9.5
      • 0710- Refill and take data with 54.5 e+/17 e- (data start at 0712)
      • 0720-0735 Refill and take data 51.7 e+/27.4 e- (data start at 0720)
        • Stop digital readout; files :
          • test_6c_cont/2005_06_29/0656/flm_paw_ind_2005_06_29_0656.dat
          • test_6c_cont/2005_06_29/0656/flm_paw_lum_2005_06_29_0656.dat
        • At 07:25:29 restart digital readout with readout now in 9x5 mode as opposed to 9x6 (this solves the problem mentioned earlier regarding the geometric data file).
        • 0727 Fix e- synch light mirrors
        • Stop data acquisition and get digital files 0725:
          • test_6c_cont/2005_06_29/0725/flm_paw_ind_2005_06_29_0725.dat
          • test_6c_cont/2005_06_29/0725/flm_paw_lum_2005_06_29_0725.dat
      • 0735-0746 Continue e- filling and take data with 47.6 e+/39.0 e-/80.7 tot (data start at 0738)
        • Restart digital data acquisition (file 0736)
      • 0747- Dump beam. Fill 1 train e+ and all trains e-
        • Files stored:
          • test_6c_cont/2005_06_29/0736/flm_paw_ind_2005_06_29_0736.dat
          • test_6c_cont/2005_06_29/0736/flm_paw_lum_2005_06_29_0736.dat
        • Unable to make this work. e- in the colliding bunch literally ejected by the transition to HEP. Can't find a direction to go before they're gone. Will need some serious work to make this set of conditions work.
      • 0800- Stop for ACCESS

June 30, 2005 (E. Tanke) Data recuperated from yesterday's 0625 and 0656 files

  • The 0625 and 0656 measurement files have been filtered such as to yield data in the correct order. All counts in the new files are valid; pulseheights are valid when they are not equal to -1. The files are:
    • old: flm_paw_ind_2005_06_29_0625.dat --> new: flm_paw_indc_2005_06_29_0625.dat
    • old: flm_paw_ind_2005_06_29_0656.dat --> new: flm_paw_indc_2005_06_29_0656.dat
  • The old files have been removed from the server.

August 2005

August 11, 2005 (E. Tanke): FLM testing: writing raw counts and raw LOI to database node for each bunch

  • FLM testing done during last night with get_6c_cont writing data to files (see /nfs/cesr/temp/ts101/FLM_data/2005_08_11 ) directory and writing raw rates and raw LOI for each bunch for the sum of the 6 vertical channels to the FLM BUN VERT node. Unlike the data written to the node, the data written to the files are averages over a user defined number of data.
    • get_6c_cont program settings: Luminosity turns=10000, geometric channel turns=250, threshold=-1200, timing=1440, averages=20
  • These data are NOT calibrated and were NOT necessarily under standard HEP conditions; however, reasonable agreement exist with the CLEO end cap measurement:

  • LOI plot for 3 bunches of Train 1 :
    LOI bunch by bunch plot

  • CLEO End Cap:
    Lumi End Cap

August 12, 2005 (E. Tanke): FLM testing: writing the sums to FLM BUN VERT

  • Data taken over last night are in /nfs/cesr/temp/ts101/FLM_data/2005_08_12. The luminosity files show rates (in stead of counts). The geometric data are averaged data; this still needs further debugging. Reverted back to writing counts to the files.
  • Using a timing scan (get_pavg), it was found that T=1440 is close to the optimum timing.
  • get_6c_cont has also been writing the sum over all bunches of the raw counts and raw LOI to database node, but when looking at the data with WINLOG, fluctuations between positive and negative values were seen. It was found that contrary to the database nodes, WINLOG only accepts 16 bit data. The data written to the node have now been rescaled for this:
    • VERT BUN RAW (locations 1 + N*6, N=0,44) have rates in [10*Hz]
    • VERT BUN LOI (locations 5 + N*6, N=0,44) are in [Hz/mA]
    • SPARE 1 (location 271) has the sum of the rates over all bunches [kHz]
    • SPARE 5 (location 275) has the sum of the LOI over all bunches [Hz/mA]

August 15, 2005 (E. Tanke): FLM testing (MS)

  • Using a timing scan (get_pavg), it was found that T=1470 is the optimum timing (at 1440 the DSP occasionally sees bunches jump in the next bucket).
  • Description of MS at: http://cesrelog.lns.cornell.edu/MSLOG/cesrc/2005/cesrc_050815.html
  • Preliminary test of an FLM display (note that here for the LOI the I was the geometric mean of the total e- and e+ currents):
    FLM display

August 16, 2005 (E. Tanke, G. Codner): FLM taken out

  • As a consequence of todays seperator problems in CESR, one will have to run at higher beam energies in order to recover the vacuum. For this reason the FLM got taken out.

August 20, 2005 (E. Tanke, G. Codner): FLM reinstalled

  • The FLM was reinstalled at 3W.
  • Upon testing with beam, the vertical channels of the FLM were found to function correctly, but geometric horizontal channels 3 and 4 seem to give a lower than expected signal.
    • Upon finding this problem, the connections to the FLM were verified and found to be in the correct order and well connected.
    • The lack of signal are seen both on the analog and digital display
    • According to the display, correct HV is present on all channels

August 22, 2005 (E. Tanke): FLM testing (MS)

  • Using HBLIMING 2 , we moved the photons towards the centre H channels; it appears indeed that channels 3 and 4 are dead. Came back with HBLIMING 2 to original setting.
  • Testing of XSHOW display. Now the L/I plots for the individual bunches are calculated with the individual bunch currents. One may also observe data from the horizontal channels, which still needs work in terms of scaling factors.
    FLM XSHOW display

August 23, 2005 (G. Codner, J. Dobbins, E. Tanke): FLM debugging (see August 20, 2005)

  • Inspected the FLM:
    • HV was found to be functional on all H channels (measured for each channel directly after the resistor on the main FLM board).
    • Disconnected each of the H channel lemos on the DSP side, one by one, and looked with a scope to check for signal (with HV on one should be able to observe spurious signal from cosmic background). Geometric channels 3 and 4 were not giving any signal, all the other connected H ones did give signal.
    • Switched off the HV and measured the resistance at the end of each lemo cable on the DSP side; one should see 10 kOhm (termination on the FLM side); did indeed see this for all of the geometric H channels.
  • The most likely possibility is radiation damage of the optical epoxy joints (which was expected to become an issue sometime in the near future - expected detector radiation damage lifetime of ~1 year) along with stresses during removal/installation, resulting in breakage.

August 24, 2005 (E. Tanke): FLM testing (MS)

  • Upon request, added L/(I*I) to the display shown above (22Aug2005). Also added the CLEO L and L/I to this display.
  • Testing of H (geometric) plots and data by connecting them in software (in get_6c_cont) to the 6 V channels. Typically V luminosity data is taken over 10k turns, geometric data over 250 turns. By applying the correct scaling one can check if rates for the 10k turns and 250 turns yield similar results. This was found to be so.
  • Two MCA scans were made, using get_24c_mca with NBIN=250, MCAMAX=500, NLOOPS=100, timing=1470) :
    • test_24c_mca/2005_08_24/FLM_PAW_MCA_2005_08_28_1257.DAT (MCAMIN= -19500)
    • test_24c_mca/2005_08_24/FLM_PAW_MCA_2005_08_28_1323.DAT (MCAMIN= -14500)
  • Data files produced today by get_6c_cont are on the server at
    • /nfs/cesr/temp/ts101/FLM_data/2005_08_24

August 25, 2005 (G. Codner, E. Tanke): FLM taken out

  • As a consequence of todays seperator problems (again) in CESR, one will have to run at higher beam energies in order to recover the vacuum. For this reason the FLM got taken out.

August 26, 2005 (M. Cole, J. Dobbins, E. Tanke, M. Ray): FLM channel testing

  • After removing the cover off of the horizontal channels, the pair of H channels 13 and 14 as well as the H pair 3 and 4 were very carefully tested for mechanical rigidity; no obvious problem was noticed (although this inspection was somewhat hampered by the fact that the pairs were left taped up and left in their mechanical fixtures. The cover was reinstalled.
  • Subsequently, the FLM was taken to 4W and the HV cables hooked up. Hooked up a scope to each of the FLM lemos on the PC board (one at a time) and looked at cosmic background; it was found that all channels, including H channels 3 and 4, were functional. Typical signals one may observe are 2-10 mV with a pulse width of about 10 ns.
  • Next, looked at cosmic ray signals by connecting the signal cables to the FLM and hooking up the other end to the scope (after disconnecting from the DSP). All H and V channels worked.
  • In order to get a higher signal rate, one may open up some of the tape on the scintillator and use a pulsed light source. However, in order to avoid mechanical contact/movement, it was decided to use a cobalt 60 source (#143, < 2 mr/h on contact) instead. Rates of several Hz or more were consistently observed.

September 2005

September 1, 2005 (G. Codner, E. Tanke): FLM reinstalled

  • The FLM was reinstalled at 3W; all channels seem functional.

September 2, 2005 (E. Tanke): FLM data

  • Data files produced last night and today by get_6c_cont are on the server at
    • /nfs/cesr/temp/ts101/FLM_data/2005_09_01
    • /nfs/cesr/temp/ts101/FLM_data/2005_09_02

September 8, 2005 (E. Tanke, M. Palmer): FLM diagnostic and calibration work

  • 12:11 Started FLM data logger program. File series 33 being written to LOG3$DISK:[CESR.FLM.DLOG]
  • SBP puts in "standard" fill which we will use for digital diagnostics and allow CLEO to run simultaneously
  • BSM alignment checkout
    • Made some mirror adjustments to get vertical electron size
    • Note that there is no VBUMPING 23 in saveset
    • Suspect that synch light may be striking burned portion of primary mirror so attempt to adjust.
    • Observe strong luminosity dependence on VB23. Adjust VNOSEING 1 and tunes to recover partially.
    • Do not observer a major change in beamsize measurement with this adjustment so perhaps we are not being badly clipped.
  • 12:15-13:30 Work on diagnosis of digital readout, in particular timing and bunch pattern issues
    • Diagnose miscellaneous issues
  • 13:35 No beam
    • 13:37 - Pedestals in POSCSR - no digital program running
      2293.3 2257.3 2830.1 2470.0 2680.0 2249.8 2440.0 2489.9
      1245.6 1115.1 1328.5 1609.2 1355.2 1190.0 1099.7 1369.4
      1146.7 1858.7 1594.4 1700.0 1417.3 1250.0 1040.0 1050.7
    • 13:42 - Pedestals in POSCSR - digital program running
      2171.8 2135.9 2718.9 2361.5 2571.5 2133.6 2323.8 2370.6
      1199.9 1069.5 1283.3 1563.5 1313.9 1145.4 1054.6 1321.9
      1098.6 1810.9 1548.8 1657.0 1373.2 1207.7 996.0 1007.0
    • 13:44 - Pedestals in HEPCSR - digital program running
      2165.6 2130.0 2714.3 2355.3 2567.1 2128.0 2318.3 2364.4
      1197.7 1066.9 1281.2 1561.2 1310.7 1143.8 1052.5 1318.8
      1096.4 1806.7 1545.6 1653.3 1369.2 1204.7 994.2 1004.9
    • 13:46 - Take MCA pedestal data (using get_24c_mca.exe) with NBIN=250, MCAMIN=-9500, MCAMAX=500, NLOOPS=100, timing=1470) :
      • File test_24c_mca/2005_09_08/1346/FLM_PAW_MCA_2005_09_08_1346.DAT

  • 13:55 - Refill
    • Restore previous pedestal and gain table for analog readout
    • Tried bunch-by-bunch readout executable with timings 1350 and 1470. With timing setting 1470 see reasonable bunch pattern agreement between bunches with current and luminosity. Note that with timing 1350, T1B1 would correspond to DSP bunch number 150 in stead of 151 ; this was verified by a timing scan which also showed that timing needs to be in the range 1470 to 1770. Further note that DSP bunch 150 is not selected in the get_6c_cont code, in which T1B1 corresponds to DSP bunch 151.

  • Recommended timing for the FLM is 1500

  • 17:00 - Take MCA data with beam (22 mA e-, 46 mA e+) with NBIN=250, MCAMIN=-9500, MCAMAX=500, NLOOPS=100, timing=1500) :
    • File test_24c_mca/2005_09_08/1700/FLM_PAW_MCA_2005_09_08_1700.DAT

September 12, 2005 (M. Palmer, E. Tanke): FLM testing

  • MCA scan (get_24c_mca) under HEP conditions with NBIN=250, MCAMIN=-9500, MCAMAX=500, NLOOPS=100, timing=1500 :
    • test_24c_mca/2005_09_12/1625/FLM_PAW_MCA_2005_09_12_1625.DAT (e+=67mA, e-=38mA)
    • test_24c_mca/2005_09_12/1805/FLM_PAW_MCA_2005_09_12_1805.DAT (e+= 0mA, e-= 0mA)

September 14, 2005 (E. Tanke): FLM testing

November 2005

November 23, 2005 (E. Tanke): Inspection of FLM

  • After the most recent CLEO run, the FLM was taken out of the machine. The scintillators and lightpipes of the 2 horizontal pairs that had been giving us intermittent problems were inspected. No visible damaged was found (one of the types of damage looked for was that due to radiation). Following this inspection, the 2 H pairs have been wrapped again and all H pairs were re-installed. The V pairs have not been touched.
  • In parallel to the above activity, a spare set of lightpipes/scintillators is being prepared.

November 28, 2005 (M. Palmer, J. Barely): FLM HV rack diagnostics

  • After finding anomalous offsets and readback problems of the FLM HV rack during Nov. 23 tests, worked on diagnosing the problems this morning
    • Find that with only the vertical HV supply (supply #1 turned on), high voltage becomes active at a command of approximately 25 V. Readbacks through the "TLM HV CTL" node appear functional and give the correct voltage.
    • Activating all 3 HV supplies causes the offset before turn-on to shift by several 10's of volts and readback no longer works.
    • After inspection, discover that the HV rack key switch disconnects the primary ground to the HV supplies' control and readback circuitry when operating with the ready-chain test relay in place. We need to review the wiring scheme with Mike Comfort to understand this fully.
  • Verify that the HV rack works properly with the ready-chain test relay in place and the rack key turned to its extreme counter-clockwise position. At this point ready for detector testing.

November 29, 2005 (M. Cole, E. Tanke): Testing of FLM in Blue Room

  • Yesterday and today testing of channels by looking at cosmic background signals. With the HV PS in the Blue Room, controlling the HV may be done using command files in the [CESR.FLM.BLUE_ROOM] directory. @HVTEST_ON 4000 5000 will cause the H channels to be set to 400 V and the V channels to 500 V. @HVTEST_OFF switches them all off.
    • Started at 400V and verified that there is no light hitting the scintillators. At this voltage, it typically takes several seconds before one sees a pulse if proper light insulation was installed. If there are light leaks, one will see pulses far more frequently.
    • Scanned up to 700 V in steps of 50 V; whenever a lightleak was found it was fixed.
    • All channels found to be well wrapped.

November 30, 2005 (E. Tanke): Status of spare FLM lightpipes/scintillators

  • 6 V sets (set= pipe + scintillator) have been polished and are ready for glueing; also some extra scintillators have been polished. 6 sets make up a complete spare for the vertical plane.
  • 6 H sets (out of 14 needed) have been polished and are ready for glueing. Also some extra scintillators have been polished.

December 2005

December 1, 2005 (M. Palmer, E. Tanke): Checkout of FLM with radiaoactive source

  • Using a cobalt 60 source, the FLM channels were checked. Rates from different channels may be different (e.g. due to differences in light coupling, scintillator to light guide and light guide to PMT, and differences in individual PMT sensitivities). The trigger level on the scope was adjusted such as to find rates that are comparable between channels. The HV was set to -600V on all H and V channels. Results (by cable number) are in the tables below.

V2 V3 V4 V5 V6 V7
Trigger level (mV) -20 -17 -18 -13 -10 -13

H1 H2 H3 H4 H6 H7 H8 H9 H10 H11 H12 H13 H14 H15
Trigger level (mV) -24 -14 -12.5 -14 -16 -12.5 -22 -8 -14 -16 -14 -13 -10 -29

  • With the HV set to 1000 CU, the voltages on the terminals were measured (see tables below).

V2 V3 V4 V5 V6 V7
HV (V) -99.5 -99.0 -98.8 -98.7 -98.9 -99.4

H1 H2 H3 H4 H6 H7 H8 H9 H10 H11 H12 H13 H14 H15
HV (V) -98.3 -98.1 -98.0 -98.2 -98.2 -98.3 -98.8 -98.1 -97.7 -97.7 -97.8 -97.8 -97.7 -97.8

  • With the HV set to 5000 CU, V2 yielded -498 V and V5 -496 V.

December 2, 2005 (E. Tanke): FLM ready for re-installation

  • Placed final wrapping on FLM lightpipes. All channels were checked once more for cosmic ray signals; all were found to be functional. FLM is ready for re-installation.

December 6, 2005 (M. Palmer, M. Cole, G. Codner, E. Tanke): FLM installed

  • FLM has been installed in CESR. The PS inside the interlock box had to be replaced as no DC voltage was present at its outputs. Proper functioning of computer control and readback of HV was verified. Cosmic ray signals were found for all channels.

December 8, 2005 (M. Palmer, E. Tanke): FLM timing scan and MCA measurement

  • Starting conditions: HEP with e- = 33 mA, e+ = 54 mA colliding
  • Timing scan with get_pavg yielded peak at 1400 (compare to 1500 on 9-Sep-2005)
  • MCA scan with get_24c_mca, T=1400, MCAMIN=-9500, MCAMAX=500, NBINS=250, NLOOPS=100. Data stored on server in
    • cesr/ts101/flm03w/test_24c_mca/2005_12_08/FLM_PAW_MCA_2005_12_08_1602.DAT

December 9, 2005 (E. Tanke): FLM MCA measurement

  • MCA scan with get_24c_mca, T=1400, MCAMIN=-50000, MCAMAX=0, NBINS=250, NLOOPS=100. Data stored on server in
    • cesr/ts101/flm03w/test_24c_mca/2005_12_09/FLM_PAW_MCA_2005_12_09_0804.DAT

December 10, 2005 (M.Palmer, E. Tanke): FLM calibrations

  • 18:02 MCA scan with get_24c_mca, T=1400, MCAMIN=-50000, MCAMAX=1000, NBINS=250, NLOOPS=100. Data stored on server in subdirectories of ~tanke/cesr/ts101/flm03w/test_24c_mca/2005_12_10
    • 1802/FLM_PAW_MCA_2005_12_10_1802.DAT Used base set of delta voltages on HV such as to correct for net gain differences between channels.
    • 1829/FLM_PAW_MCA_2005_12_10_1829.DAT Trippled the delta voltages of the 1802 measurement.
    • 1949/FLM_PAW_MCA_2005_12_10_1949.DAT Used nominal voltages (no offsets), but T=1500.
  • Started test_6c_cont; this program loads the digital FLM database nodes; TH=-1200, T=1500, #averages=25, #turns=10000; #turns (individual channels)=250

December 11, 2005 (M.Palmer, E. Tanke): FLM calibrations (continued)

  • MCA scans with get_24c_mca, T=1500, MCAMIN=-50000, MCAMAX=1000, NBINS=250, NLOOPS=100. Data stored on server in subdirectories of ~tanke/cesr/ts101/flm03w/test_24c_mca/2005_12_11
  • 1312/FLM_PAW_MCA_2005_12_11_1312.DAT Used voltages on HV:
    • Voltage set by [cesr.flm.com]adjust_hv.com;14 with v_offset=100,h_offset=150,scale=4.
    • Furthermore d1=0,d2=-10,d3=-4,d4=-6,d5=4,d6=4,d7=0,d8=0,
    • d9=-8,d10=4,d11=6,d12=4,d13=0,d14=0,d15=6,d16 =-8,
    • d17=10,d18=5,d19=2,d20=5,d21=4,d22=8,d23=-16 and d24=0.
    • Note: hv_vert=7250+ v_offset and hv_horz=7500+h_offset.
    • Read back of voltages in table below:

Vertical Horizontal
Channel HV(V) Rate HV(V) Rate
1 -762.5 230 -796.5 60
2 -702.0 673 -810.0 86
3 -761.5 3030 -752.0 218
4 -730.0 8205 -787.5 809
5 -745.5 3361 -807.0 1828
6 -719.5 598 -788.0 1870
7 * * -809.0 552
8 * * -803.5 173
9 * * -799.5 88
10 * * -776.0 72
11 * * -785.0 44
12 * * -706.0 31
13 * * -750.5 23
14 * * -791.5 21
VSUM= 16097 HSUM= 5875
VLoI= 3792 HLoI= 1384

  • 1326/FLM_PAW_MCA_2005_12_11_1326.DAT Horizontally beam moved to higher H channels and vertically beam moved to lower V channels.
  • Summed data for the 1312 and 1326 measurements can be found in ~tanke/cesr/ts101/flm03w/test_24c_mca/summeddata/2005_12_11

December 12, 2005 (M.Palmer, E. Tanke): FLM calibrations (continued)

December 15, 2005 (M.Palmer, E. Tanke): FLM calibrations (continued)

  • New set of voltages for HV applied (see table below)
Vertical Horizontal
Channel HV(V) HV(V)
1 -762.5 -772.0
2 -686.0 -809.5
3 -762.0 -732.0
4 -722.0 -787.0
5 -745.0 -802.5
6 -719.0 -771.5
7 * -798.5
8 * -811.5
9 * -795.0
10 * -772.0
11 * -776.5
12 * -698.0
13 * -730.0
14 * -792.0

  • MCA scan with settings like 11-Dec-2005. Data file: 2005_12_15/0722/FLM_PAW_MCA_2005_12_15_0722.DAT
  • MCA scan with faster DSP code (modified MCA routine: dsp_mca_fast.c). This faster version can be loaded from the control system by typing @flm_dlf from the test_24c_mca directory. dsp_mca_fast works without pedestals and the number of bins is hardwired to 250. Data file: 2005_12_15/0747/FLM_PAW_MCA_2005_12_15_0747.DAT

December 17, 2005 (M.Palmer, E. Tanke): FLM plotting programs

  • Beta versions of 3 FLM plotting programs, which may be of particular use during tuning, have been written. The time plots automatically refresh at the end of the default time span; at this point data are shifted and the plot subsequently continues with the last half of the previous time span worth of data still visible. The following 3 executables are in the [CESR.PALMER.TS101.PLOTPROGS] directory:
    • FLM_BUNCHES_TUNING (contains histogram and time plots related to luminosity, both global and bunch-by-bunch)
    • FLM_BUNCHES_LOI_FAST (plots the L/I for individual bunches as a function of time; update rate approximately once per second, time span 5 minutes)
    • FLM_BUNCHES_LOI_SLOW (plots the L/I for individual bunches as a function of time; update rate approximately once per 12 seconds, user defined time span (up to 12 hours)

December 19, 2005 (M.Palmer, E. Tanke): FLM calibrations (continued)

  • MCA scan using the original (slow) get_24c_mca with settings like 11-Dec-2005 except for the number of loops. Data files:
    • 2005_12_19/0934/FLM_PAW_MCA_2005_12_19_0934.DAT (500 loops; electron loss occured while recording)
    • 2005_12_19/1005/FLM_PAW_MCA_2005_12_19_1005.DAT (500 loops)
    • 2005_12_19/1040/FLM_PAW_MCA_2005_12_19_1040.DAT (200 loops; exciting different channels)
    • 2005_12_19/1057/FLM_PAW_MCA_2005_12_19_1057.DAT (200 loops; exciting different channels)

December 21, 2005 (C. Strohman, J. Dobbins, E. Tanke): Preliminary testing of FLMA

  • A first series of tests have been conducted on the new FLMA module, which will include accumulator boards.

December 23, 2005 (M. Palmer, E. Tanke): Installation of new FLM HV settings

  • VOLTAGE CORRECTIONS (in magnitude - NOT SIGNED)
HV Ch Dig Ch Vmag Adjust
1 7 0
2 6 5.6
3 5 -2.4
4 4 8
5 3 8
6 2 4
7 1 1.6
8 0 0
9 23 4
10 22 1.6
11 21 4.8
12 20 -4
13 19 0
14 18 -5.6
15 17 8.8
16 16 1.6
17 15 9.6
18 14 4.8
19 13 -4
20 12 4
21 11 6.4
22 10 8
23 9 -8
24 8 0

  • Deltas installed into [cesr.flm.com]adjust_hv4.com
  • Baseline voltages: V = -725 V H = -750 V
HV Ch Volts
1 -0
2 -684.6
3 -708.6
4 -719.0
5 -759.0
6 -755.0
7 -736.6
8 -0
9 -717.0
10 -778.6
11 -789.8
12 -753.0
13 -0
14 -751.4
15 -787.8
16 -714.6
17 -814.6
18 -773.8
19 -765.0
20 -789.0
21 -787.4
22 -813.0
23 -685.0
24 -0

December 27, 2005 (M.Palmer, E. Tanke): FLM calibrations (continued)

  • MCA scan using the original (slow) get_24c_mca Data files:
    • 2005_12_27/1026/FLM_PAW_MCA_2005_12_27_1026.DAT (500 loops; T=1500, NBINS=250, MCAMIN=-20000, MCAMAX=1000)
    • 2005_12_27/1110/FLM_PAW_MCA_2005_12_27_1110.DAT (500 loops; T=1500, NBINS=250, MCAMIN=-10000, MCAMAX=1000, V channels only)

December 28, 2005 (M.Palmer, E. Tanke): FLM calibrations (continued)

  • New voltages applied to PMT HV - HV values based on 12/27/05 1026 calibration
$! Modify style so that di's are in 10ths of volts
$ v_offset = 100
$ h_offset = 150
$ hv_vert = 7250 + v_offset
$ hv_horz = 7500 + h_offset
$ scale = 1
$ conv  = 1 * scale
$ d1   =   0
$ d2   = -569
$ d3   = -270
$ d4   = -290
$ d5   =  157
$ d6   =  206
$ d7   =   19
$ d8   =   0
$ d9   = -546
$ d10  =  125
$ d11  =  229
$ d12  = -189
$ d13  =   0
$ d14  = -181
$ d15  =  228
$ d16  = -527
$ d17  =  452
$ d18  =   66
$ d19  =  -49
$ d20  =  220
$ d21  =  218
$ d22  =  478
$ d23  = -825
$ d24  =   0
    • Observed operating voltages and rates after the above calibration installed.
+----------------------------------------------------------+
|          |       Vertical        |      Horizontal       |
|  Channel |   HV(V)  |    Rate    |   HV(V)  |    Rate    |
+----------------------------------------------------------+
      1       -762.5        556       -761.5         98
      2       -685.0       1607       -815.5        131
      3       -767.0       8996       -731.5        245
      4       -719.0       9679       -793.5        807
      5       -747.5       3242       -806.0       1919
      6       -714.0        687       -774.5       2655
      7                               -808.0       1620
      8                               -820.0        620
      9                               -796.0        233
     10                               -764.0        131
     11                               -766.5         78
     12                               -687.5         56
     13                               -727.5         41
     14                               -794.0         36
+----------------------------------------------------------+

      VSUM =        24767            HSUM =  8670

************************************************************
  • MCA scan using the original (slow) get_24c_mca Data files:
    • 2005_12_28/0934/FLM_PAW_MCA_2005_12_28_0934.DAT (500 loops; T=1500, NBINS=250, MCAMIN=-20000, MCAMAX=1000)
    • Recalibrated
    • New voltages applied to PMT HV
    • New deltas are:
$ d1   =   0
$ d2   = -579
$ d3   = -270
$ d4   = -352
$ d5   =  102
$ d6   =  209
$ d7   =   21
$ d8   =   0
$ d9   = -563
$ d10  =  119
$ d11  =  219
$ d12  = -204
$ d13  =   0
$ d14  = -204
$ d15  =  226
$ d16  = -530
$ d17  =  437
$ d18  =   66
$ d19  =  -77
$ d20  =  220
$ d21  =  220
$ d22  =  480
$ d23  = -826
$ d24  =   0
  • New voltages and rates are:
+----------------------------------------------------------+
|          |       Vertical        |      Horizontal       |
|  Channel |   HV(V)  |    Rate    |   HV(V)  |    Rate    |
+----------------------------------------------------------+
      1       -757.5        551       -760.5        113
      2       -683.5       1430       -814.0        156
      3       -767.0       7888       -731.0        261
      4       -719.0      10981       -793.5        837
      5       -747.5       4505       -804.5       1955
      6       -708.5        851       -775.0       2801
      7                               -808.0       1709
      8                               -820.0        668
      9                               -795.5        248
     10                               -760.5        137
     11                               -764.5         80
     12                               -687.5         61
     13                               -725.5         45
     14                               -794.0         40
+----------------------------------------------------------+

      VSUM =        26206            HSUM =  9111
  • New MCA calibration:
    • 2005_12_28/1146/FLM_PAW_MCA_2005_12_28_1146.DAT (500 loops; T=1500, NBINS=250, MCAMIN=-20000, MCAMAX=1000)
    • This calibration run to verify operating point
    • Further changes to deltas no larger than 14 in magnitude. This is essentially the noise level so we will leave things as is.

December 29, 2005 (E. Tanke): FLM calibrations (continued)

  • MCA scan using the original (slow) get_24c_mca; repeat of yesterday's 1146 measurement to check stability. Data file:
    • 2005_12_29/1121/FLM_PAW_MCA_2005_12_29_1121.DAT (500 loops; T=1500, NBINS=250, MCAMIN=-20000, MCAMAX=1000)

December 31, 2005 (C. Strohman, M. Palmer, E. Tanke): Testing of new FLMA with accumulator board installed on one ADC board (V channels)

  • This week a series of basic tests have been conducted on the accumulator board of the new FLMA, including:
    • Accumulator card registers for indirect addressing
    • Bunch pattern registers
    • Hit Mask register
    • Pedestal and Threshold registers
    • Rate Period registers (Bunch, Fast, Slow, Geo)
    • Rate Data registers (Bunch, Fast, Slow, Geo Global, Geo Channel)
  • Furthermore, a look up table was loaded and tested ('hi hit' channels only); note that the channel mapping in place will be changed as it currently runs from digital channel 0..5 (it should run from geo channel 1..6).
    • Tests performed by connecting one or several channels with a test signal on 1 bunch and reading over 10000 turns:
      Digi Channel(s)     Fast Global
         connected           Rate
             0              10000
             1              10000
             2              10000
             3              10000
             4              10000
             5              10000
             6                0
             7                0
            0,1             10000
            0,2             20000
            1,2             10000
            1,3             20000
           0,1,2            20000
           1,2,3            20000
          0,1,2,3           30000
          1,2,3,4           30000
           0,2,4            30000
           1,3,5            30000
  • Slow Global Rates in the above tests were identical to the Fast Global Rates
Topic revision: r62 - 30 Oct 2007, DevinBougie
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding CLASSE Wiki? Send feedback